# MO-YOLO: End-to-End Multiple-Object Tracking Method with YOLO and Decoder

### v3

@@ 4.1 Datasets and Metrics
This paper introduces MO-YOLO, an innovative MOT method proficient in addressing intricate scenarios involving occlusions, interactions, and diverse motions. To validate its performance, we evaluate the algorithm on three demanding datasets: DanceTrack[29], MOT17[24], and KITTI[12].
@@ 4.2 Implementation Details
The rationale for adopting MO-YOLO originates from our observation of the high training costs and suboptimal inference speed associated with MOTR. Thus, our aim is to propose a new baseline model surpassing MOTR’s performance. To ensure a fair comparison, we exclusively employ methods enhancing tracking performance, while omitting improvement techniques utilized in MeMOTR [10], MOTRv2 [47], MOTRv3 [40], and CO-MOT [38]. To achieve a comparable parameter count with MOTR, we simultaneously utilize YOLOv8-L [14] and RT-DETR-L [21] in constructing MO-YOLO (henceforth referred to as MO-YOLO-L concerning model size). The initial weights are obtained from a model pre-trained on the COCO dataset [18].
@@
During the initial training, batch sizes were set to 32 on the 4090 and 16 on the 2080ti. Subsequent stages utilized a uniform batch size of 1. For MOT17, like MOTR, we integrated extra data from the CrowdedHuman dataset [28], employing random displacement to generate video segments with pseudo-trajectories. Training comprised 120 epochs in the first stage, followed by 30 epochs in the second. The third stage extended to 20 epochs with TBSP, or 55 without. For Dancetrack, ensuring parity without extra data like MOTRv2, the first stage was 15 epochs, the second was 6, and the third, if TBSP was used, was 6 epochs, or 15 without TBSP. For KITTI, initial training involved 80 epochs, followed by 25 in the second stage. The third stage consisted of 30 epochs with TBSP or 70 without. Additionally, approximately 5k images from BDD100k [41] were selected for additional training.
Note: With the exception of Table 4(b), all data concerning MO-YOLO in this paper pertains to MO-YOLO-L. Additionally, unless stated otherwise, the provided data pertains to the final end-to-end model.
@@ 4.3 Performance Evaluation and Comparison
In Table 1, a comprehensive comparison of training time and speed was conducted between MO-YOLO and the MOTR series.
@@
MO-YOLO stage1 1 2080ti 12.1 hours
MO-YOLO stage2 1 2080ti 30.8 hours
MO-YOLO(TBSP) 1 2080ti 43.4 hours
MO-YOLO 1 2080ti 68.7 hours
@@
MO-YOLO(TBSP) 18.5
MO-YOLO 19.6
Table 1: Comparison of training time (including all previous stages) and speed between MO-YOLO and MOTR series on a single V100 GPU. The data for the MOTR series is sourced from [42], [40] or their open source address.
It is crucial to note that the datasets used for comparing training time and inference speed exhibit variations. This disparity primarily arises from the act that the publicly disclosed training times and inference speeds of the MOTR series are based on different datasets.
MO-YOLO demonstrates a substantial advantage over the MOTR series, particularly in terms of training time and speed. The primary reason behind this efficiency gain lies in the architectural simplicity of MO-YOLO. Unlike MOTR, MO-YOLO adopts a streamlined structure that eliminates the need for an encoder, resulting in a more computationally frugal design. This reduction in complexity enables MO-YOLO to achieve training completion within 68.7 hours, utilizing only 1 Nvidia GeForce 2080ti GPU. In contrast, MOTR relies on 8 Nvidia GeForce 2080ti GPUs and takes about 96 hours for training. The absence of an encoder in MO-YOLO contributes to its swift training process, emphasizing the model’s efficiency in resource utilization.
On the Dancetrack test dataset, MO-YOLO(TBSP) achieves a speed of 18.5 frames per second (FPS) using a single V100, while the MOTR series attains speeds of 9.5 FPS (MOTR), 6.9 FPS (MOTRv2), and 10.6 FPS (MOTRv3). Particularly noteworthy is the remarkable 19.6 FPS achieved by MO-YOLO, emphasizing its outstanding performance in inference speed.
Table 2: Performance comparison between MO-YOLO and existing methods on the Dancetrack[29] dataset under the private detection protocols. Numbers are highlighted in bold when they correspond to either MO-YOLO or the superior metric from MOTR.
Methods HOTA AssA DetA IDF1 MOTA
@@
ByteTrack[45] 47.3 31.3 71.6 52.5 89.5
GTR[49] 48.0 31.9 72.5 50.3 84.7
DST-track[6] 51.9 34.6 82.3 51.0 84.9
OC-SORT[5] 54.6 38.0 80.4 54.2 89.4
@@
MeMOTR[10] 63.4 52.3 77.0 65.5 85.4
MOTR[42] 54.2 40.2 73.5 51.5 79.7
MO-YOLO(Ours) 52.0 35.7 75.9 51.4 83.5
Table 2, Table 3, and Table 4 exhibit the final results of end-to-end MO-YOLO on the Dancetrack, MOT17, and KITTI datasets, respectively. In contrast to MOTR, MO-YOLO demonstrates competitive performance across multiple metrics on both datasets. It is important to highlight that the performance of MeMOTR is provided in Tables 2 and 3. The inclusion of MeMOTR’s performance aims to underscore the considerable potential for advancement and enhancement in both MOTR and our MO-YOLO. MeMOTR[10] integrates additional modules and performance-enhancing techniques, along with supplementary datasets during training. Consequently, comparing its tracking performance directly with that of MO-YOLO and MOTR may not be entirely fair.
Table 3: Performance comparison between MO-YOLO and existing methods on the MOT17[24] dataset.
method HOTA AssA DetA IDF1 MOTA
@@
FairMOT[46] 59.3 58.0 60.9 72.3 73.7
ByteTrack[45] 62.8 62.2 63.8 77.2 78.9
StrongSORT[49] 63.5 63.7 63.6 78.3 78.5
UTM[39] 64.0 62.5 65.9 78.7 81.8
OC-SORT[5] 63.2 63.2 63.2 77.5 78.0
Deep OC-SORT[22] 64.9 65.9 64.1 80.6 79.4
@@
MeMOTR[10] 58.8 58.4 59.6 69.0 72.5
MOTR[42] 57.2 55.8 58.9 59.4 71.9
MO-YOLO(Ours) 55.5 55.9 55.8 70.1 66.8
For Dancetrack (Table 2), MO-YOLO achieves comparable results to MOTR, with slightly lower scores in HOTA and AssA but outperforming MOTR in DetA. Specifically, MO-YOLO achieves an impressive MOTA, indicating its overall superior performance in multi-object tracking tasks.
In the MOT17 (Table 3), MO-YOLO demonstrates competitive performance, particularly in handling complex tracking scenarios, with competitive scores in HOTA and AssA. Despite trailing slightly behind MOTR in these metrics, MO-YOLO compensates with a comparable DetA and excels in IDF1 by almost 10 points. This advantage can be attributed to MO-YOLO’s simplified network structure, which enables competitive tracking accuracy with reduced training data demand. Additionally, MO-YOLO maintains robust overall performance, as evidenced by its high MOTA score, highlighting its suitability for real-world tracking applications.
Table 4: Performance comparison between MO-YOLO and existing methods on the KITTI[12] dataset.
Car Pedestrian
@@
CenterTr[48] 73.0 88.8 71.1 40.3 53.8 36.9
PermaTr[30] 77.4 90.8 77.6 47.4 65.0 43.6
PolarMOT[15] 75.1 80.0 76.9 43.6 46.9 48.1
TripletTrack[23] 73.6 84.3 74.7 50.0 46.5 57.8
@@
MO-YOLO 72.2 83.2 73.8 51.5 56.8 58.4
Table 4 provides an overview of MO-YOLO’s performance on the KITTI dataset. MO-YOLO achieves competitive results, with an HOTA of 70.2 and 82.9 for Car tracking and Pedestrian tracking, respectively. These results position MO-YOLO as a strong contender in multi-object tracking on the KITTI.
In conclusion, MO-YOLO stands out as a highly advantageous model in the realm of multi-object tracking, demonstrating superior object detection accuracy and overall tracking performance compared to MOTR, as evidenced by competitive results on diverse datasets, including Dancetrack, MOT17, and KITTI. The notable improvements observed in MeMOTR relative to MOTR suggest that MO-YOLO has significant untapped potential for performance enhancement. Positioned as a robust alternative, MO-YOLO can serve as an excellent new baseline, driving advancements in end-to-end models within the field of multi-object tracking.
@@ 4.4 Ablation Study
To assess the proposed training strategy and the efficacy of TBSP, ablation experiments were conducted on the Dancetrack dataset. The training set includes the entire Dancetrack training data, with testing involving 25 sequences from the validation set. Additionally, the impact of two methods for generating Query Positions—utilizing the formula from DAB-DETR[19] or employing an MLP—on the model’s performance was investigated. Notably, MO-YOLO-L, MO-YOLO-M, and MO-YOLO-S differ based on their foundations in YOLOv8-L,YOLOv8-M, and YOLOv8-S, respectively, providing insights into model variations and their implications.
(a) Performance of MO-YOLO at different training stages Training Stage HOTA AssA MOTA Initial 38.5 21.6 76.8 Intermediate 44.5 25.5 84.2 Final (TBSP) 51.9 34.6 84.9 Final 53.6 35.9 84.7 (b) Performance across different model sizes Model HOTA MOTA FPS MO-YOLO-L 53.6 84.7 19.6 MO-YOLO-M 49.7 82.9 22.7 MO-YOLO-S 42.9 79.3 26.2
(c) The impact of $\delta_{iou}$ on model performance when using TBSP $\delta_{iou}$ HOTA AssA MOTA 0.3 49.7 32.3 79.1 0.5 51.9 34.6 84.9 0.7 51.3 34.8 83.2 0.9 50.0 32.8 81.3 (d) Effect of query position generation strategy Strategy HOTA MOTA FPS MLP 53.3 85.1 18.5 Formula 53.6 84.7 19.6
Table 5: Ablation Studies on Our Proposed MO-YOLO on the DanceTrack validation set.
The performance of MO-YOLO improves consistently across different training stages, peaking in the final stage. Different model sizes (MO-YOLO-L, MO-YOLO-M, MO-YOLO-S) achieve a balance between accuracy and computational efficiency. The similarity in their speeds is likely attributable to the consistent scale of the decoder. The choice of $\delta_{iou}$ impacts precision and recall trade-offs, with higher values leading to a slight decrease in performance. Position generation strategies, MLP and Formula, yield similar performance in terms of detection accuracy, with slight differences in FPS. Overall, the ablation studies demonstrate the effectiveness of MO-YOLO across different training conditions and model configurations, showcasing its adaptability and versatility in MOT field.
Figure 5: MO-YOLO and MOTR failed in the same scenario. The probable cause may lie in the insufficient scale of the MOT17 dataset.
@@ 5 Conclusion and Limitation
This paper presents MO-YOLO, an efficient and lightweight end-to-end multi-object tracking model that optimizes computational resources. By integrating YOLO and RT-DETR principles, the study establishes a novel tracking paradigm with accelerated training strategies. While MO-YOLO outperforms MOTR in resource efficiency, it marginally lags behind in performance, attributed to factors like YOLO’s detection capabilities and reduced input image dimensions ($640\times 640$). Despite enhancing computational efficiency, the dimension reduction raises concerns about potential information loss, impacting overall effectiveness. Additionally, latent issues from MOTR may surface in MO-YOLO, as depicted in Fig.5. This research contributes to advancing real-time computer vision applications.
Appendix 0.A Appendix
0.A.1 Disclaimer
The data used in this paper is sourced from publicly available datasets related to human subjects. While every effort has been made to ensure the accuracy and reliability of the data, the authors make no representations or warranties of any kind, express or implied, about the completeness, accuracy, reliability, suitability, or availability with respect to the data. The authors shall not be liable for any loss or damage, including but not limited to indirect or consequential loss or damage, arising from the use of the data in this research paper. Users of the data are encouraged to independently verify the accuracy and reliability of the data before making any decisions or taking any actions based on it.
0.A.2 More Comprehensive Analysis of Experimental Results and Model Parameter
The initial motivation for building MO-YOLO was not because of its slow inference speed, but rather because of its high training resource requirements. With 8 V100 GPUs using pre-trained weights, it took approximately 84 hours to train 200 epochs on an extended MOT17[24] with only about 10,000 images. Each epoch took around 0.42 hours, and on average, each GPU could train on approximately 3,000 images per hour.
It is worth noting that MOTR[42] is based on Deformable-DETR[50], which was trained on the MSCOCO dataset[18] with approximately 110,000 images. In single-scale mode (which can be considered similar to MOTR’s training), using eight V100 GPUs, each epoch took approximately 0.4 hours, and on average, each GPU could train on around 30,000 images per hour.
Upon careful consideration of the reasons, the first aspect to be contemplated is the batch size. The collective average loss (CAL) mechanism and the tracklet-aware label assignment (TALA) strategy employed in MOTR enable the network to learn some latent temporal information, consequently restricting the batch size on a single card to 1. Setting the batch size to 1 may result in underutilization of the corresponding computational resources (e.g., GPU, TPU, or other possible devices). Additionally, MOTR is trained in a one-shot manner, requiring the model not only to grasp motion information but also to learn object appearance features. Hence, we proposed a multi-step training strategy to first allow the model to learn appearance features before proceeding to learn object motion characteristics. Experimental results confirm the effectiveness of this strategy.
As for the TBSP, its initial purpose was simply to provide the detection network with a basic tracking capability. However, it also pleasantly surprised us by further reducing training time. Here, we also provide the required rounds and results if TALA were to be used directly without the TBSP strategy, as shown in Table 6. The training here is conducted on a single NVIDIA 4090 GPU. And this training process is not included in the first stage
@@ 0.A.2 More Comprehensive Analysis of Experimental Results and Model Parameter
Table 6: Comparison of MO-YOLO with and without using TBSP on the Dancetrack val dataset.
Initial Intermediate Final HOTA MOTA Epochs Total Time
53.6 84.7 27 56.7 hours
52.9 83.8 40 85.2 hours
In order to provide a more comprehensive illustration of the characteristics of our model, we list the sizes of different models, namely MO-YOLO-L (45.8M), MO-YOLO-M (29.9M), and MO-YOLO-S (19.7M), as depicted in Fig. 6.
Figure 6: The model sizes of MOTR and MO-YOLO, as well as their FPS and HOTA on the Dancetrack.
The table compares the performance of MO-YOLO with and without using the TBSP strategy on the Dancetrack validation dataset across different training phases: Initial, Intermediate, and Final. Comparison of the two rows reveals that employing TBSP throughout all phases leads to slightly better performance in terms of HOTA and MOTA, with a significantly lower total training time. This underscores the importance of the TBSP strategy in reducing the overall training duration while maintaining or even improving tracking accuracy.
We hypothesize that the efficacy of the TBSP strategy may stem from its endowment of the model with a degree of autonomy, facilitating the preliminary acquisition of object motion information that would otherwise necessitate numerous training iterations. Analogously, the strategy resembles the use of training wheels in bicycle riding, which assist novices in achieving smooth operation while providing an opportunity to experience the sensation of riding. Consequently, this approach aids in the comprehensive mastery of the skill. We posit that a comparable scenario may manifest within this context.
0.A.3 RT-MOTR
At the same time, to demonstrate the advantages of our MO-YOLO models constructed using YOLO and RT-DETR relative to end-to-end tracking models built directly on RT-DETR, we have also constructed such an end-to-end tracking model based on RT-DETR (referred to as RT-MOTR), and conducted performance testing on it.
NOTE: in Table 7, both the train time and epochs refer to the second and third stages like Table 6. For the first stage, both models were trained for 15 epochs, and the training device is a single 4090, while the testing device is a single V100.
The table compares the performance of MO-YOLO and RT-MOTR on the Dancetrack validation dataset. It presents various metrics including HOTA (Higher Order Tracking Accuracy), training epochs, training speed (hours per epoch), total training time, and frames per second (FPS).
MO-YOLO achieves a HOTA of 53.6 with 27 training epochs, an average training speed of 2.08 hours per epoch, a total training time of 56.7 hours, and an FPS of 19.6.
In comparison, RT-MOTR achieves a HOTA of 53.4 with 35 training epochs, an average training speed of 2.27 hours per epoch, a total training time of 90.8 hours, and an FPS of 20.3.
Comparing the performance metrics of the two methods, we observe that MO-YOLO slightly outperforms RT-MOTR in terms of HOTA, while also having a shorter training time albeit with a slightly lower FPS. This indicates that MO-YOLO exhibits advantages in both tracking accuracy and training efficiency.
Table 7: Comparison of MO-YOLO and RT-MOTR on the Dancetrack val dataset.
Methods HOTA Train Epochs Speed (hours/epoch) Train Times FPS
MO-YOLO 53.6 27 2.08 56.7 h 19.6
RT-MOTR 53.4 35 2.27 79.5 h 20.3
Regarding why the speeds of these two models are comparable, our experimental findings suggest that the primary time-consuming component is the decoder responsible for tracking. Therefore, optimizing the decoder represents a crucial area for future investigation.
This outcome underscores the rationale behind integrating elements from YOLOv8 and RT-DETR to construct a rapid end-to-end tracking network. Furthermore, the observed disparity in training efficiency between MO-YOLO and RT-DETR can be attributed to the former’s notably simpler architecture.
Moreover, our decision to adopt YOLO as the backbone of the network is motivated by its propensity for swifter updates compared to DETR-based models. Recent developments such as YOLOv9[34] exemplify this trend. While YOLOv9 may trade off training time for enhanced precision and speed (resulting in at least a twofold increase in training duration compared to YOLOv7[33]), we anticipate future iterations will offer refined improvements. Such advancements will, in turn, facilitate further enhancements to MO-YOLO.

### v4

@@ 4.1 Datasets and Metrics
To validate DecoderTracker’s performance, we evaluate the algorithm on three demanding datasets: DanceTrack[32], MOT17[33], and KITTI[34].
@@ 4.2 Implementation Details
The rationale for adopting DecoderTracker originates from our observation of the high training costs and suboptimal inference speed associated with MOTR. Thus, our aim is to propose a new baseline model surpassing MOTR’s performance. To ensure a fair comparison, we exclusively employ methods enhancing tracking performance, while omitting improvement techniques utilized in MeMOTR [10], MOTRv2 [18], MOTRv3 [9], and CO-MOT [8].
@@
During the initial training, batch sizes were set to 32 on the 4090 and 16 on the 2080ti. Subsequent stages utilized a uniform batch size of 1. For MOT17, like MOTR, we integrated extra data from the CrowdedHuman [37], employing random displacement to generate video segments with pseudo-trajectories. Training comprised 120 epochs in the first stage, followed by 30 epochs in the second. The third stage extended to 20 epochs with TBSP, or 55 without. For Dancetrack, ensuring parity without extra data like MOTRv2, the first stage was 15 epochs, the second was 9, and the third, if TBSP was used, was 9 epochs, or 18 without TBSP. For KITTI, initial training involved 80 epochs, followed by 25 in the second stage. The third stage consisted of 30 epochs with TBSP or 70 without. Additionally, approximately 5k images from BDD100k [38] were selected for additional training.
@@ 4.3 Performance Evaluation and Comparison
In Table 1, a comprehensive comparison of training time and speed was conducted between DecoderTracker and the MOTR series. It is crucial to note that the datasets used for comparing training time and inference speed exhibit variations. This disparity primarily arises from the act that the publicly disclosed training times and inference speeds of the MOTR series are based on different datasets.
@@
DecoderTracker stage1 1 2080ti 12.1 hours
DecoderTracker stage2 1 2080ti 30.8 hours
DecoderTracker(TBSP) 1 2080ti 43.4 hours
DecoderTracker 1 2080ti 68.7 hours
@@
DecoderTracker(TBSP) 18.5
DecoderTracker 19.6
Table 1: Comparison of training time (including all previous stages) and speed between DecoderTracker and MOTR series on a single V100 GPU. The data for the MOTR series is sourced from [3], [9] or their open source address.
DecoderTracker demonstrates a substantial advantage over the MOTR series, particularly in terms of training time and speed. The primary reason behind this efficiency gain lies in the architectural simplicity of DecoderTracker. Unlike MOTR, DecoderTracker adopts a streamlined structure that eliminates the need for an encoder, resulting in a more computationally frugal design. This reduction in complexity enables DecoderTracker to achieve training completion within 68.7 hours, utilizing only 1 Nvidia GeForce 2080ti GPU. In contrast, MOTR relies on 8 Nvidia GeForce 2080ti GPUs and takes about 96 hours for training. The absence of an encoder in DecoderTracker contributes to its swift training process, emphasizing the model’s efficiency in resource utilization.
Table 2: Performance comparison between DecoderTracker and existing methods on the Dancetrack[32] benchmark. Numbers are highlighted in bold when they correspond to either DecoderTracker or the superior metric from MOTR.
Methods HOTA IDF1 DetA AssA MOTA
@@
ByteTrack[14] 47.3 52.5 71.6 31.3 89.5
GTR[39] 48.0 50.3 72.5 31.9 84.7
DST-track[40] 51.9 51.0 82.3 34.6 84.9
OC-SORT[41] 54.1 54.2 80.4 38.0 89.4
@@
MeMOTR[10] 63.4 65.5 77.0 52.3 85.4
MOTR[3] 54.2 51.5 73.5 40.2 87.6
DecoderTracker(Ours) 54.2 54.2 78.7 37.4 87.8
On the Dancetrack test benchmark, DecoderTracker(TBSP) achieves a speed of 18.5 frames per second (FPS) using a single V100, while the MOTR series attains speeds of 9.5 FPS (MOTR), 6.9 FPS (MOTRv2), and 10.6 FPS (MOTRv3). Particularly noteworthy is the remarkable 19.6 FPS achieved by DecoderTracker, emphasizing its outstanding performance in inference speed. Table 2, Table 3, and Table 4 exhibit the final results of end-to-end DecoderTracker on the Dancetrack, MOT17, and KITTI benchmark, respectively. In contrast to MOTR, DecoderTracker demonstrates competitive performance across multiple metrics on both datasets. It is important to highlight that the performance of MeMOTR is provided in Tables 2 and 3. The inclusion of MeMOTR’s performance aims to underscore the considerable potential for advancement and enhancement in both MOTR and our DecoderTracker. MeMOTR[10] integrates additional modules and performance-enhancing techniques, along with supplementary datasets during training. Consequently, comparing its tracking performance directly with that of DecoderTracker and MOTR may not be entirely fair.
Table 3: Performance comparison between DecoderTracker and existing methods on the MOT17[33] set.
method HOTA IDF1 DetA AssA MOTA
@@
FairMOT[42] 59.3 72.3 60.9 58.0 73.7
ByteTrack[14] 62.8 77.2 63.8 62.2 78.9
StrongSORT[39] 63.5 78.3 63.6 63.7 78.5
UTM[43] 64.0 78.7 65.9 62.5 81.8
OC-SORT[41] 63.2 77.5 63.2 63.2 78.0
Deep OC-SORT[6] 64.9 80.6 64.1 65.9 79.4
@@
MeMOTR[10] 58.8 69.0 59.6 58.4 72.5
MOTR[3] 57.2 59.4 58.9 55.8 71.9
DecoderTracker(Ours) 55.7 70.5 55.7 56.0 66.5
For Dancetrack (Table 2), DecoderTracker achieves comparable results to MOTR, with same scores in HOTA but outperforming MOTR in DetA. Specifically, DecoderTracker achieves an impressive MOTA, indicating its overall superior performance in multi-object tracking tasks.
In the MOT17 (Table 3), DecoderTracker demonstrates competitive performance, particularly in handling complex tracking scenarios, with competitive scores in HOTA and AssA. Despite trailing slightly behind MOTR in these metrics, DecoderTracker compensates with a comparable DetA and excels in IDF1 by almost 10 points. This advantage can be attributed to DecoderTracker’s simplified network structure, which enables competitive tracking accuracy with reduced training data demand. Additionally, DecoderTracker maintains robust overall performance, as evidenced by its high MOTA score, highlighting its suitability for real-world tracking applications.
Table 4: Performance comparison between DecoderTracker and existing methods on the KITTI[34] set.
Pedestrian Car
@@
CenterTr[44] 40.3 53.8 36.9 73.0 88.8 71.1
PermaTr[45] 47.4 65.0 43.6 77.4 90.8 77.6
PolarMOT[46] 43.6 46.9 48.1 75.1 80.0 76.9
TripletTrack[47] 50.0 46.5 57.8 73.6 84.3 74.7
@@
DecoderTracker 51.5 56.8 58.4 72.2 83.2 73.8
Table 4 provides an overview of DecoderTracker’s performance on the KITTI benchmark. DecoderTracker achieves competitive results, with an HOTA of 70.2 and 82.9 for Car tracking and Pedestrian tracking, respectively. These results position DecoderTracker as a strong contender in MOT on the KITTI.
@@ 4.4 Ablation Study
To assess the proposed training strategy and the efficacy of TBSP, ablation experiments were conducted on the Dancetrack benchmark. The training set includes the entire Dancetrack training data, with testing involving 25 sequences from the validation set. Additionally, the impact of two methods for generating query positions—utilizing the sinusoidal functions for encodings from DAB-DETR[48] or employing an MLP, on the model’s performance was investigated.
(a) Performance of DecoderTracker at different training stages
Training Stage HOTA AssA MOTA
First 38.5 21.6 76.8
Second 44.5 25.5 84.2
Final (TBSP) 51.9 34.6 84.9
Final 53.6 35.9 84.7
(b) The presence or absence of TAN during the second training stages can significantly impact the performance of the model.
Model HOTA AssA MOTA
Second 44.5 25.5 84.2
Second(w/o) 42.1 21.4 80.8
Final 53.6 35.9 84.7
Final(w/o) 53.4 36.1 84.5
(c) The impact of $\delta_{iou}$ on model performance when using TBSP
$\delta_{iou}$ HOTA AssA MOTA
0.3 49.7 32.3 79.1
0.5 51.9 34.6 84.9
0.7 51.3 34.8 83.2
0.9 50.0 32.8 81.3
(d) Effect of query position generation strategy
Setting HOTA MOTA FPS
sin/cos 53.3 85.1 18.5
MLP 53.6 84.7 19.6
Table 5: Ablation Studies on Our Proposed DecoderTracker on the DanceTrack validation set.
The performance of DecoderTracker improves consistently across different training stages, peaking in the final stage. The choice of $\delta_{iou}$ impacts precision and recall trade-offs, with higher values leading to a slight decrease in performance. Position generation strategies, MLP and sinusoidal functions, yield similar performance in terms of detection accuracy, with slight differences in FPS. Overall, the ablation studies demonstrate the effectiveness of DecoderTracker across different training conditions and model configurations, showcasing its adaptability and versatility in MOT field.
It is noteworthy that in Table 5(b), we demonstrate the impact of the presence or absence of TAN during the second phase of training on the network’s tracking performance at different stages. It can be observed that the presence or absence of TAN in the second phase has almost no effect on the final network performance. However, the absence of TAN does impact the tracking performance during the second phase, which also validates the effectiveness of our proposed TBSP strategy in enhancing tracking capabilities.
@@ 5 Conclusion and Limitation
This paper presents an efficient decoder-only end-to-end MOT (Multi-Object Tracking) model that boasts faster inference speeds and reduced training times. Additionally, this paper introduces new training strategies aimed at accelerating model convergence. Although device limitations prevented us from validating our training strategies on the original MOTR model, we believe these strategies are effective. This research provides a more efficient benchmark model for the multi-object tracking field compared to MOTR, laying a foundation for the emergence of real-time end-to-end MOT models and offering insights for future development. In subsequent work, we should further explore more efficient network architectures to construct real-time MOT models. Of course, we also plan to incorporate techniques from MOTRv2, MOTRv3, and CO-MOT into DecoderTracker to enhance tracking performance.
Appendix A More Discusstion About TBSP
Figure 5: Refinement process for DecoderTracker on partially trained track queries. When DecoderTracker undergoes comprehensive training with the TALA strategy, TBSP will no longer be utilized. At that point, the track query generation method of DecoderTracker aligns with MOTR, as illustrated in the Fig. 1.
@@ Appendix A More Discusstion About TBSP
Table 6: Comparison of DecoderTracker with and without using TBSP on the Dancetrack val set.
Setting HOTA MOTA Epochs Total Time
TBSP 53.6 84.7 27 56.7 hours
w/o 53.4 84.2 40 85.2 hours
The initial reason for developing the MO-YOLO model was not due to slow inference speed, but rather the extensive training resources required. Using eight V100 GPUs with pretrained weights, training on an extended MOT17 dataset, which contained approximately 10,000 images, required about 84 hours for 200 epochs. Each epoch took around 0.42 hours, with each GPU processing about 3,000 images per hour.
Fortunately, our DecoderTracker model has addressed this issue. Although its inference speed is not yet real-time, its convergence rate is already satisfactory. Although we have not achieved real-time inference speeds, we have significantly improved the speed of MOTR.
Regarding the TBSP, its original purpose was simply to endow the detection network with basic tracking functionality. However, it also provided some unexpected benefits, such as further reducing training time. Below, we list the results and required iterations if training was conducted directly using TALA, as shown in Table 6.
We believe the effectiveness of the TBSP strategy may be due to its function as a form of weak or self-supervised training for motion features, compared to TALA. Many examples demonstrate that undergoing weakly supervised training before strongly supervised training can enhance performance and speed up convergence. For instance, in the NLP field, large language models like BERT[49] and GPT[1] initially train on a vast amount of unlabeled or weakly labeled data to learn basic language patterns. Subsequently, these models undergo strong supervision fine-tuning for specific tasks like sentiment analysis or question answering. This approach has been empirically proven to significantly boost the performance of LLMs.
In our field of computer vision, this technique is also quite common, with the most classic example being the SimCLR series[50, 51]. Typically, these models are pretrained on unlabeled data to learn image features, then fine-tuned on strongly supervised tasks such as classification or detection.
We believe that the TBSP, a weak supervision process, enables the model to preliminarily learn object motion information that would otherwise require many more training iterations. It’s akin to learning to ride a bicycle where training wheels help beginners to ride without obstacles while also giving them a feel for biking, which aids in eventually mastering the skill. A similar situation may exist here.
Although TBSP did not enhance model performance, it reduced the number of iterations needed for convergence, which might be considered its greatest contribution.
Appendix B Design of IMFENet
The design of this feature network is intended to enhance the model’s efficiency and increase its inference speed. Thus, our design primarily references the network structure of YOLOv8. The backbone structure of our model is almost identical to that of YOLOv8. Our main modification involves changes to the neck structure in YOLOv8, to ensure that its output aligns with the input requirements of the decoder.
Furthermore, we have made some modifications to the C2F module in YOLOv8, renaming our modified module to C2M. This change is specifically reflected in the BottleNeckF module. In YOLOv8, there is a BottleNeck module that differs slightly between the neck and backbone; we eliminated this difference and replaced the 3x3 convolution in the middle with a 1x1 convolution to further reduce the computational load. These modifications are illustrated in Figure 6, and their impact on performance is shown in Table 7.
Figure 6: The difference between our proposed model and YOLOv8 in terms of modules. Our modules are more lightweight. Specific model details can refer to the figure.3.
Figure 7: DecoderTracker and MOTR failed in the same scenario. The probable cause may lie in the insufficient scale of the MOT17 benchmark.
Table 7: The impact of our network and YOLOv8 structure on DecoderTracker performance on the Dancetrack val set. Training time does not include the first stage
Setting HOTA IDF1 MOTA AssA FPS
IMFENet 53.6 53.8 84.7 35.9 19.6
YOLOv8 52.4 51.4 84.2 36.1 18.9
Appendix C Limitation
Undoubtedly, the greatest limitation of DecoderTracker, similar to that of MOTR, is its performance on the MOT17 benchmark, which does not match that of tracking-by-detection models. Both models fail under similar circumstances, as illustrated in Figure.7. This issue may, like with MOTR, stem from the insufficient data volume on the MOT17 benchmark, which does not adequately train the model. Clearly, despite this limitation, DecoderTracker still represents a significant improvement over MOTR.